# MMedAgent-RL: Optimizing Multi-Agent Collaboration for Multimodal Medical Reasoning

### v2

@@ Abstract
Medical Large Vision-Language Models (Med-LVLMs) have shown strong potential in multimodal diagnostic tasks. However, existing single-agent models struggle to generalize across diverse medical specialties, limiting their performance. Recent efforts introduce multi-agent collaboration frameworks inspired by clinical workflows, where general practitioners (GPs) and specialists interact in a fixed sequence. Despite improvements, these static pipelines lack flexibility and adaptability in reasoning. To address this, we propose MMedAgent-RL, a reinforcement learning (RL)-based multi-agent framework that enables dynamic, optimized collaboration among medical agents. Specifically, we train two GP agents based on Qwen2.5-VL via RL: the triage doctor learns to assign patients to appropriate specialties, while the attending physician integrates the judgments from multi-specialists and its own knowledge to make final decisions. To address the inconsistency in specialist outputs, we introduce a curriculum learning (CL)-guided RL strategy that progressively teaches the attending physician to balance between imitating specialists and correcting their mistakes. Experiments on five medical VQA benchmarks demonstrate that MMedAgent-RL not only outperforms both open-source and proprietary Med-LVLMs, but also exhibits human-like reasoning patterns. Notably, it achieves an average performance gain of 20.7% over supervised fine-tuning baselines.
@@ 1 Introduction
Large Vision-Language Models (LVLMs) are becoming increasingly proficient in visual understanding and reasoning [30, 31, 64, 2, 8]. This advancement is also making a significant impact in the biomedical domain, where Medical Large Vision-Language Models (Med-LVLMs) have demonstrated great potential in enabling intelligent diagnostic applications [24, 36, 38, 56, 6]. However, as shown in Figure 1 (a) left, although a single Med-LVLM can be trained with a large amount of data and show promise results to some extent, it is challenging for a single model to handle diagnostic expertise from different subfields (e.g., radiology, pathology, etc.).
Therefore, some recent works propose using multi-agent collaboration [55, 26, 16, 14] to solve medical tasks [27, 20, 48], where different models act as specialists or general practitioners, collaborating and discussing to arrive at a final answer, improving overall performance compared to a single agent. These works follow the steps of simulating a hospital visit process and adopt a General Practitioner (GP) $\rightarrow$ Specialist $\rightarrow$ GP workflow. First, the general practitioner (i.e., the triage doctor) classifies the patient based on the consultation questions and images and selects the appropriate department from several predefined specialties. Then, specialist doctors from the relevant departments provide their diagnoses. Finally, the general practitioner (i.e., attending physician) makes the final decision based on the images, consultation questions, and the diagnostic results from multiple specialists. However, as illustrated in Figure 1 (a) middle, such workflows are inherently static. Such interaction pattern between agents is fixed and predetermined, which limits the system’s capacity for flexible, optimized reasoning across multiple modalities.
@@
To address this challenge, motivated by the success of Reinforcement Learning (RL) [47]-driven reasoning [19, 13, 50], as shown in Figure 1 (a) right, we perform a stage-wise training on two GPs based on Qwen2.5-VL [3] via reinforcement learning, namely the triage doctor and the attending physician. Specifically, first, for the first GP, i.e., triage doctor, we utilize the image modality information provided by the dataset itself to reinforce the triage doctor, such as pathology slides $\rightarrow$ Pathologist, ensuring that the triage doctor can accurately assign patients to the appropriate department. Then, we use powerful proprietary models like GPT-4o [39] to play the role of the specialist doctors and generate initial judgments. Finally, the second GP, i.e., attending physician, integrates domain knowledge from multiple specialists and their own judgment to make the final decision. Here, during the process of the general practitioner integrating specialist doctor information, while specialist doctors provide valuable domain knowledge, their judgments are not always perfectly accurate. These inconsistencies in specialist performance can introduce noise into model training, preventing the model from simply memorizing or blindly replicating their outputs. Instead, the model must learn to generalize beyond potentially flawed expert judgments. To address this, inspired by Curriculum Learning (CL) [4, 42, 10], which enables models to be trained progressively on increasingly difficult tasks, we implement a reinforcement learning approach based on CL, aiming to help the model gradually learn to leverage the knowledge of specialist doctors. This stepwise learning facilitates better knowledge integration and application by first teaching the model to simply imitate the answers of specialist doctors, and then gradually guiding it to reason out the correct answers when the specialists make mistakes. Specifically, we use the accuracy of specialist results as a flag to classify the training data by difficulty: specialist results that are completely correct are labeled as easy, partially correct as medium, and completely incorrect as hard. In this way, we design a three-stage curriculum reinforcement learning process for optimizing the attending physician to handle diverse specialist results, including how to accurately reference specialist knowledge and when to rely on its own understanding to solve problems.
@@
The primary contribution of this paper is MMedAgent-RL, an RL-driven framework optimized for multi-agent collaboration in improving medical reasoning. Empirical results on five medical multimodal datasets, shows that the model performs exceptionally well not only on in-domain datasets but also on out-of-domain datasets, outperforming a series of both open-source and proprietary LVLMs, exceeding SFT method [40] by 20.7%. In addition to the quantitative performance, we also observed that the model’s reasoning process resembles human doctors’ thought processes: it first defines the disease mentioned in the question, carefully analyzes the image, and then determines if it can fully match the disease definition.
Figure 2: Overview of MMedAgent-RL, a RL-driven multi-agent framework designed to enhance the multimodal medical reasoning. It simulates the clinical loop of General Practitioner (GP) $\rightarrow$ Specialists $\rightarrow$ GP. MMedAgent-RL uses GRPO [13] to optimize the triage doctor (the first GP) in order to improve triage accuracy. Then, powerful proprietary LVLMs are used as the specialist doctors for the assigned department. Finally, curriculum learning [4, 42] and RL are combined to progressively train the attending physician (the second GP), who integrates the diverse opinions of specialists and makes robust decisions under varying levels of expert reliability.
@@ 2 Preliminaries
Group Relative Policy Optimization (GRPO). Group Relative Policy Optimization (GRPO) [13] is a reinforcement learning method that avoids training a critic by using intra-group relative rewards to optimize the policy. For each query $x$, the model samples $G$ responses $\{y^{(1)},\ldots,y^{(G)}\}$, which are scored to get rewards $\{R_{1},\ldots,R_{G}\}$. GRPO computes normalized advantages and updates the policy with a PPO-style clipped objective [44]:
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{x,\{y_{i}\}}\Bigg{[}\frac{1}{G}\sum_{i=1}^{G}\Bigg{(}\min\Big{(}r_{i}A_{i},\,\text{clip}(r_{i},1-\epsilon,1+\epsilon)A_{i}\Big{)}-\beta\,\mathbb{D}_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\Bigg{)}\Bigg{]},r_{i}=\frac{\pi_{\theta}(y_{i}\mid x)}{\pi_{\text{old}}(y_{i}\mid x)},$ (1)
@@
Fine-tuning with GRPO. We fine-tune our model $\pi_{\theta}$ using GRPO to enhance the reasoning process. Each training sample $x\in\mathcal{D}$ consists of an image $x_{v}$, a textual prompt $x_{t}$ composed of a user query and system message, and a reference output format containing <think> and <answer> tags. Given input $x=(x_{v},x_{t})$, the model samples a group of candidate outputs $\{o^{(1)},\ldots,o^{(G)}\}\sim\pi_{\theta}(y\mid x)$, where each $o^{(i)}$ includes both reasoning trace and final answer. A reward function scores each output based on correctness of the final answer $r_{\text{accuracy}}$ and formatting compliance with XML-style tags $r_{\text{format}}$. Relative scores $\{R_{i}\}$ are computed and normalized to obtain advantages $\{A_{i}\}$, which are used to update $\pi_{\theta}$ following the GRPO objective (Eq. (1)).
@@ 3 Methodology
In this section, as illustrated in Figure 2, we will present MMedAgent-RL, a novel RL-driven multi-agent framework for multimodal medical reasoning by emulating a structured clinical workflow. Our approach begins with the first General Practitioner (GP) leveraging the input information to intelligently select the most appropriate medical department for further consultation. To optimize the accuracy of this initial triage decision, we employ GRPO [13] to train and refine the triage doctor’s capabilities. Subsequently, the case is referred to a panel of specialist doctors, each represented by a powerful proprietary LVLM specialized in the identified department. These LVLMs analyze the provided multimodal data and generate expert opinions relevant to their respective fields. Finally, the process culminates with the second GP, acting as the attending physician, who integrates the diverse insights from the specialist LVLMs along with their own medical knowledge to arrive at a comprehensive and robust final judgment. We will delve into the specifics of each stage as follows:
@@ 3.1 Initial Departmental Triage
In real-world medical treatment processes, the first critical step when a patient visits a hospital is triage [1], which involves assigning the patient to the appropriate department based on their symptoms. Previous works simulated hospital visits using LLMs or LVLMs, with Agent Hospital [27] employing an LLM for triage, and MDAgents [20] enhancing this by dynamically selecting collaboration structures and using GPT-4o for department assignment. However, these methods rely on a rigid and predefined approach for department assignment as they cannot update the model based on new data.
To address this challenge, in our approach, the first step is to optimize the general practitioner $a_{\text{GP}}^{\text{triage}}$ who acts as the triage doctor (i.e., policy $\pi_{\theta_{\text{GP}}^{\text{triage}}}$), using Qwen2.5-VL [3] as the base model. Here, we use the image modality information provided by the dataset itself as ground truth labels $y^{*}$ to train the triage model. For example, pathology slides $\rightarrow$ pathologist (e.g., PathVQA contains pathology slide images and is thus assigned to pathologists), chest X-ray images $\rightarrow$ radiologist (e.g., VQA-RAD includes various types of radiological images and is therefore assigned to radiologists), ensuring that the triage model can accurately assign patients to the appropriate medical specialty.
Specifically, when prompting the triage doctor, we provide $k$ candidate specialties. In our setup, $k$ is set to 7, including Pathologist, Radiologist, Surgeon, Oncologist, Endocrinologist, Ophthalmologist, and Dermatologist, which broadly cover the main departments involved in the data. Our aim is not only to improve triage accuracy, but also to strengthen the model’s reasoning process, helping explain why a particular triage recommendation is made. Thus, we use GRPO [13] as the base RL algorithm. At this stage, the reward function adopts a rule-based format with rewards $R_{\text{format}}\in\{0,0.5\}$ and accuracy rewards $R_{\text{accuracy}}\in\{0,1\}$. A reward is given when the output format meets the required criteria and the chosen specialty is correct. This stage can optimize the triage doctor’s performance, improving their ability to select the appropriate specialty $\small d=\arg\max_{k}\pi_{\theta_{\text{GP}}^{\text{triage}}}(k\mid x)$, and lays a foundation for subsequently acting as the corresponding specialist.
@@ 3.2 Role-Playing Specialists Offer Valuable Insights
After obtaining the department from the triage doctor, following previous work using LLMs or LVLMs for medical discussions [27, 48, 23, 49], we utilize several powerful models as specialist doctors $\small a_{\text{SP}}^{(d)}$ to provide relatively accurate preliminary judgments. This facilitates subsequent reference by the attending physician. In our setup, we use responses from $e$ specialists as references for each sample. We only require the specialist doctors to independently provide expert opinions $\small y_{d}\sim\pi_{\theta_{\text{SP}}^{(d)}}(y\mid x)$ within their specialty, without engaging in complex interactions. This ensures system efficiency and avoids majority voting that could overshadow minority opinions, leaving the final decision to the attending physician.
@@ 3.3 Evolving Decisions by Attending Physician via Ongoing Collaboration
Figure 3: (a) The average performance of both general practitioners (GP) and specialists is below 70%, which suggests a misalignment issue in multi-agent collaboration. Over-reliance on specialists’ opinions or unilateral decisions by GPs can both lead to suboptimal outcomes. (b) Since specialists perform inconsistently across different cases, this poses a challenge for GPs when making decisions. Using all data for reinforcement fine-tuning can easily trap the model in a locally suboptimal solution (left). In contrast, our C-MARL approach enables the model to progressively accomplish sub-goals in a three-stage process and ultimately reach a globally optimal solution (right).
@@
After getting the responses from the specialists, we then integrate their knowledge into the final general practitioner designed to support the final diagnostic decision. The final decision-making agent, namely the attending physician, plays the most crucial role throughout the diagnostic process, as they must synthesize diverse expert opinions and draw upon their own clinical expertise to arrive at a final judgment. This poses significant challenges for the attending physician, as the specialists’ conclusions are not always fully reliable. For example, as shown in Figure 3 (a), although specialists generally outperform general practitioners, their overall accuracy is still below 70%. Therefore, over-reliance on specialist input can lead to suboptimal outcomes. Secondly, different specialists may offer conflicting interpretations of the same case, creating misalignment issues. If the model is unable to reconcile its internal reasoning with external expert input, it risks compounding errors. For instance, while majority voting may help mitigate the influence of less competent specialists, it can also suppress minority views—including potentially the only correct one. As illustrated in Figure 3 (a), such multi-agent collaboration (GPT-4o $\rightarrow$ Qwen2.5-VL or GPT-4o) can yield adverse effects when the model is not properly aligned with the nature and limitations of expert knowledge.
To address these challenges, as shown in Figure 3 (b), we draw inspiration from curriculum learning [4, 42, 37], which emphasizes the importance of organizing learning experiences in a meaningful progression, i.e., from easier to harder tasks. Motivated by this principle, we propose the Curriculum-based Reinforcement Learning tailored for Multi-Agent Collaboration (C-MARL), which gradually increases task difficulty during training. This stepwise learning approach promotes better knowledge integration and application by first training the model to mimic specialist doctors’ answers, and then gradually encouraging it to reason independently and correct errors when specialists make mistakes.
@@
Specifically, unlike previous curriculum learning approaches that define difficulty based on problem formulation or data domains, we categorize tasks based on the accuracy of specialists’ diagnoses $\small y_{d}\sim\pi_{\theta_{\text{SP}}^{(d)}}(y_{d}\mid x)$, denoted by $s=\text{Acc}(y_{d},y^{*})$. The dataset is divided into three levels: fully correct specialist results ($s=1$) are labeled as easy, partially correct results ($10<s<$ 1) are labeled as medium, and completely incorrect results ($s=0$) are labeled as hard. The datasets corresponding to the three levels are denoted as $\mathcal{D}_{\text{easy}},\mathcal{D}_{\text{medium}},\mathcal{D}_{\text{hard}}$, respectively, and $\mathcal{D}=\mathcal{D}_{\text{easy}}\cup\mathcal{D}_{\text{medium}}\cup\mathcal{D}_{\text{hard}}$. Based on these data of three categories, as shown in Algorithm 1, we design a three-stage curriculum reinforcement learning process to optimize the attending physician’s ability to handle different types of specialist knowledge, such as when to accurately leverage specialist knowledge and when to rely on their own understanding to solve problems. Specifically, at each stage, we adopt GRPO as the base RL algorithm. For each query $x$, the attending physician generates a group of $G$ responses $\{y_{\text{final}}^{(1)},y_{\text{final}}^{(2)},...,y_{\text{final}}^{(G)}\}$, where each $y_{\text{final}}\sim\pi_{\theta_{\text{GP}}^{\text{attend}}}(y\mid x,y_{d})$. Subsequently, for each query with $G$ responses, GRPO computes the relative advantage of each response based on their rewards, which are determined by a format reward $R_{\text{format}}\in\{0,0.5\}$ and an accuracy reward $R_{\text{accuracy}}\in\{0,1\}$, avoiding training a complex critic model. The relative advantage is computed as $\small A_{i}=\frac{R_{i}-\text{mean}(\{R_{j}\}_{j=0}^{G})}{\text{std}(\{R_{j}\}_{j=0}^{G})}$. Then, as illustrated in Equation 1, GRPO uses the computed advantages to optimize the policy via the PPO-clip loss function, directly incorporating a KL penalty term to improve model performance. Notably, we apply different initial KL coefficients for tasks of varying difficulty levels, ensuring the model can gradually adapt to scenarios reflecting the performance of multiple specialist doctors.
@@
Input: Task dataset $\mathcal{D}=\{x_{v}^{(i)},x_{t}^{(i)},y^{*(i)}\}_{i=1}^{N}$, policy model $\pi_{\theta}$, old policy $\pi_{\text{old}}$, group size $G$, responses from specialists $y_{d}^{(i)}$.

### v3

@@ Abstract
Medical Large Vision-Language Models (Med-LVLMs) have shown strong potential in multimodal diagnostic tasks. However, existing single-agent models struggle to generalize across diverse medical specialties, limiting their performance. Recent efforts introduce multi-agent collaboration frameworks inspired by clinical workflows, where general practitioners (GPs) and specialists interact in a fixed sequence. Despite improvements, these static pipelines lack flexibility and adaptability in reasoning. To address this, we propose MMedAgent-RL, a reinforcement learning (RL)-based multi-agent framework that enables dynamic, optimized collaboration among medical agents. Specifically, we train two GP agents based on Qwen2.5-VL via RL: the triage doctor learns to assign patients to appropriate specialties, while the attending physician integrates the judgments from multi-specialists and its own knowledge to make final decisions. To address the inconsistency in specialist outputs, we introduce a curriculum learning (CL)-guided RL strategy with dynamic entropy regulation, progressively teaching the attending physician to balance between imitating specialists and correcting their mistakes. Experiments on five medical VQA benchmarks demonstrate that MMedAgent-RL outperforms both open-source and proprietary Med-LVLMs. Notably, it achieves an average performance gain of 23.6% over strong baselines.
@@ 1 Introduction
Large Vision-Language Models (LVLMs) are becoming increasingly proficient in visual understanding and reasoning (Liu et al., 2024a; b; Zhu et al., 2023; Bai et al., 2023; Chen et al., 2024c). This advancement is also making a significant impact in the biomedical domain, where Medical Large Vision-Language Models (Med-LVLMs) have demonstrated great potential in enabling intelligent diagnostic applications (Li et al., 2023; Moor et al., 2023; Nath et al., 2024). However, as shown in Figure 1 (a) left, although a single Med-LVLM can be trained with a large amount of data and show promise results to some extent, it is challenging for a single model to handle diagnostic expertise from different subfields (e.g., radiology, pathology, etc.).
Therefore, some recent works propose using multi-agent collaboration to solve medical tasks (Li et al., 2024c; Kim et al., 2024; Tang et al., 2024), where different models act as specialists or general practitioners, collaborating and discussing to arrive at a final answer, improving overall performance compared to a single agent. These works follow the steps of simulating a hospital visit process and adopt a General Practitioner (GP) $\rightarrow$ Specialist $\rightarrow$ GP workflow. First, the general practitioner (i.e., the triage doctor) classifies the patient based on the consultation questions and images and selects the appropriate department from several predefined specialties. Then, specialist doctors from the relevant departments provide their diagnoses. Finally, the general practitioner (i.e., attending physician) makes the final decision based on the images, consultation questions, and the diagnostic results from multiple specialists. However, as illustrated in Figure 1 (a) middle, such workflows are inherently static. Such interaction pattern between agents is fixed and predetermined, which limits the system’s capacity for flexible, optimized reasoning across multiple modalities.
To address this challenge, motivated by the success of Reinforcement Learning (RL)-driven reasoning (Jaech et al., 2024; Guo et al., 2025; Team et al., 2025), as shown in Figure 1 (a) right, we perform a stage-wise training on two GPs based on Qwen2.5-VL (Bai et al., 2025) via reinforcement learning, namely the triage doctor and the attending physician.
@@
Specifically, first, for the first GP, i.e., triage doctor, we utilize the image modality information provided by the dataset itself to reinforce the triage doctor, such as pathology slides $\rightarrow$ Pathologist, ensuring that the triage doctor can accurately assign patients to the appropriate department. Then, we use powerful proprietary models like GPT (OpenAI, 2025) to play the role of the specialist doctors and generate initial judgments. Finally, the second GP, i.e., attending physician, integrates domain knowledge from multiple specialists and their own judgment to make the final decision. Here, during the process of the general practitioner integrating specialist doctor information, while specialist doctors provide valuable domain knowledge, their judgments are not always perfectly accurate. These inconsistencies in specialist performance can introduce noise into model training, preventing the model from simply memorizing or blindly replicating their outputs. Instead, the model must learn to generalize beyond potentially flawed expert judgments.
To address this, inspired by Curriculum Learning (CL) (Bengio et al., 2009; Pentina et al., 2015; Deng et al., 2025), which enables models to be trained progressively on increasingly difficult tasks, we implement an entropy-aware reinforcement learning approach based on CL, aiming to help the model gradually learn to leverage the knowledge of specialist doctors. The core principle is that the attending physician agent faces a dynamic exploration-exploitation dilemma when integrating specialist opinions: it must learn when to trust and adopt a consensus and when to challenge it and search for novel diagnostic paths (Cui et al., 2025; Wang et al., 2025b). Specifically, we use the accuracy of specialist results as a flag to classify the training data by difficulty: specialist results that are completely correct are labeled as easy, partially correct as medium, and completely incorrect as hard. In this way, we design a three-stage curriculum reinforcement learning process for optimizing the attending physician to handle diverse specialist results, where the model starts with low-entropy policies to exploit reliable “easy” cases and gradually increases its policy entropy to explore and correct flawed judgments in “hard” cases.
@@
The primary contribution of this paper is MMedAgent-RL, an RL-driven framework optimized for multi-agent collaboration in improving medical reasoning. Empirical results on five medical multimodal datasets, shows that the model performs exceptionally well not only on in-domain datasets but also on out-of-domain datasets, outperforming a series of both open-source and proprietary LVLMs, exceeding SFT method by 23.6%.
@@ 2 Preliminaries
Group Relative Policy Optimization (GRPO). Group Relative Policy Optimization (GRPO) (Guo et al., 2025) is a reinforcement learning method that avoids training a critic by using intra-group relative rewards to optimize the policy. For each query $x$, the model samples $G$ responses $\{y^{(1)},\ldots,y^{(G)}\}$, which are scored to get rewards $\{R_{1},\ldots,R_{G}\}$. GRPO computes normalized advantages and updates the policy with a PPO-style clipped objective (Schulman et al., 2017):
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{x,\{y_{i}\}}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\Bigg(\min\Big(r_{i}A_{i},\,\text{clip}(r_{i},1-\epsilon,1+\epsilon)A_{i}\Big)-\beta\,\mathbb{D}_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\Bigg)\Bigg],r_{i}=\frac{\pi_{\theta}(y_{i}\mid x)}{\pi_{\text{old}}(y_{i}\mid x)},$ (2.1)
@@
Figure 2: Overview of MMedAgent-RL, a RL-driven multi-agent framework designed for multimodal medical reasoning. It simulates the clinical loop of General Practitioner (GP) $\rightarrow$ Specialists $\rightarrow$ GP. First, MMedAgent-RL optimizes the triage doctor (the first GP) to improve triage accuracy. Then, proprietary LVLMs are used as the specialist doctors for the assigned department. Finally, curriculum learning and RL are combined to progressively train the attending physician (the second GP), who integrates the specialist knowledge and makes robust decisions.
@@ 3 Methodology
In this section, as illustrated in Figure 2, we will present MMedAgent-RL, a RL-driven multi-agent framework for multimodal medical reasoning by emulating a structured clinical workflow. Our approach begins with the first General Practitioner (GP) leveraging the input to select the appropriate medical department for further consultation. To optimize the initial triage decision, we employ GRPO to refine the triage doctor’s capabilities. Then the case is referred to a panel of specialist doctors, each served by a proprietary LVLM specialized in the identified department, analyze data and provide expert opinions. Finally, the process culminates with the second GP, acting as the attending physician, integrates all specialist opinions to form a final diagnosis. We will delve into each stage as follows:
@@ 3.1 Initial Departmental Triage
In real-world medical treatment, particularly for complex diagnoses, the workflow fundamentally relies on a Multi-Disciplinary Team (MDT) process (Abo-Hamad and Arisha, 2013): moving from initial assessment and triage to expert consultation, and finally, synthesis by an attending physician. This collaborative paradigm has been validated in recent medical AI literature as an effective abstraction for simulating clinical decision-making. Previous works like Agent Hospital (Li et al., 2024c) and MDAgents (Kim et al., 2024) have adopted similar role-based structures, employing LLMs for triage or dynamic collaboration. However, these methods typically rely on rigid or predefined assignment strategies and lack the ability to update the model based on new data.
To address this challenge, the first step is to optimize the general practitioner $a_{\text{GP}}^{\text{triage}}$ who acts as the triage doctor (i.e., policy $\pi_{\theta_{\text{GP}}^{\text{triage}}}$). Here, we use the image modality information provided by the dataset itself as ground truth labels $y^{*}$ to train the triage model. For example, pathology slides $\rightarrow$ pathologist (e.g., PathVQA contains pathology slide images and is thus assigned to pathologists), ensuring that the triage model can accurately assign patients to the appropriate medical specialty.
Specifically, when prompting the triage doctor, we provide $k$ candidate specialties. In our setup, $k$ is set to 7, including Pathologist, Radiologist, Surgeon, Oncologist, Endocrinologist, Ophthalmologist, and Dermatologist, which broadly cover the main departments involved in the data. Our aim is not only to improve triage accuracy, but also to strengthen the model’s reasoning process, helping explain why a particular triage recommendation is made. Thus, we use GRPO as the base RL algorithm. At this stage, the reward function adopts a rule-based format with rewards $R_{\text{format}}\in\{0,0.5\}$ and accuracy rewards $R_{\text{accuracy}}\in\{0,1\}$. A reward is given when the output format meets the required criteria and the chosen specialty is correct. This stage can optimize the triage doctor’s performance, improving their ability to select the appropriate specialty $d=\arg\max_{k}\pi_{\theta_{\text{GP}}^{\text{triage}}}(k\mid x)$, and lays a foundation for subsequently acting as the corresponding specialist.
@@ 3.2 Role-Playing Specialists Offer Valuable Insights
After obtaining the department from the triage doctor, following previous work using LLMs or LVLMs for medical discussions (Li et al., 2024c; Tang et al., 2024; Li et al., 2024a), we utilize several powerful models as specialist doctors $a_{\text{SP}}^{(d)}$ to provide relatively accurate preliminary judgments. This facilitates subsequent reference by the attending physician. In our setup, we use responses from $e$ specialists as references for each sample. We only require the specialist doctors to independently provide expert opinions $y_{d}\sim\pi_{\theta_{\text{SP}}^{(d)}}(y\mid x)$ within their specialty, without engaging in complex interactions. This ensures system efficiency and avoids majority voting that could overshadow minority opinions, leaving the final decision to the attending physician.
@@ 3.3 Evolving Decisions by Attending Physician via Ongoing Collaboration
After getting the responses from the specialists, we then integrate their knowledge into the final general practitioner designed to support the final diagnostic decision. The final decision-making agent, namely the attending physician, plays the most crucial role throughout the diagnostic process, as they must synthesize diverse expert opinions and draw upon their own clinical expertise to arrive at a final judgment. This poses significant challenges for the attending physician, as the specialists’ conclusions are not always fully reliable. Over-reliance on specialist input can lead to suboptimal outcomes. Secondly, different specialists may offer conflicting interpretations of the same case, creating misalignment issues. If the model is unable to reconcile its internal reasoning with external expert input, it risks compounding errors. For instance, while majority voting may help mitigate the influence of less competent specialists, it can also suppress minority views, including potentially the only correct one. Such multi-agent collaboration can yield adverse effects when the model is not properly aligned with the nature and limitations of expert knowledge.
To address these challenges, we draw inspiration from curriculum learning (Bengio et al., 2009), which emphasizes the importance of organizing learning experiences in a meaningful progression, i.e., from easier to harder tasks. Motivated by this principle, see in Alg. 1, we propose the Curriculum-based Entropy-Aware Reinforcement Learning for Multi-Agent Collaboration (C-MARL). The core principle of this framework is that the attending physician agent faces a dynamic exploration-exploitation dilemma when integrating specialist opinions: it must learn when to trust and adopt a consensus (exploit) and when to challenge it and search for novel diagnostic paths (explore). We posit that the policy entropy is the central mechanism for navigating this trade-off (Cui et al., 2025). Therefore, C-MARL employs a purpose-designed curriculum to actively shape the policy entropy of the attending physician, enabling it to adapt to specialist information of varying reliability.
@@
Curriculum Design based on Specialist Reliability. Unlike previous curriculum learning approaches that define difficulty based on problem formulation or data domains, we categorize tasks based on the accuracy of specialists’ diagnoses $y_{d}\sim\pi_{\theta_{\text{SP}}^{(d)}}(y_{d}\mid x)$, denoted by $s=\text{Acc}(y_{d},y^{*})$. The dataset is divided into three levels: fully correct specialist results ($s=1$) are labeled as easy, partially correct results ($0<s<1$) are labeled as medium, and completely incorrect results ($s=0$) are labeled as hard. The datasets corresponding to the three levels are denoted as $\mathcal{D}_{\text{easy}},\mathcal{D}_{\text{medium}},\mathcal{D}_{\text{hard}}$, respectively, and $\mathcal{D}=\mathcal{D}_{\text{easy}}\cup\mathcal{D}_{\text{medium}}\cup\mathcal{D}_{\text{hard}}$. Based on these data of three categories, we design a three-stage curriculum reinforcement learning process to optimize the attending physician’s ability to handle different types of specialist knowledge, such as when to accurately leverage specialist knowledge and when to rely on their own understanding to solve problems.
Reinforcement Learning with Dynamic Entropy Regulation. We utilize GRPO as our base RL algorithm. For each query $x$, the attending physician $\pi_{\theta_{\text{GP}}^{\text{attend}}}$ generates a group of $G$ responses $\{y_{\text{final}}^{(i)}\}_{i=1}^{G}$. The reward $R_{i}$ for each response is determined by a format reward $R_{\text{format}}\in\{0,0.5\}$ and an accuracy reward $R_{\text{accuracy}}\in\{0,1\}$, and GRPO calculates the relative advantage $A_{i}=\frac{R_{i}-\text{mean}(\{R_{j}\})}{\text{std}(\{R_{j}\})}$ for the policy update. To achieve dynamic entropy control, we introduce an entropy regularization term into the standard GRPO objective function:
$\mathcal{J}_{\text{C-MARL}}(\theta)=\mathbb{E}\left[\mathcal{J}_{\text{GRPO}}(\theta)+\gamma_{s}\cdot H_{t}(\pi_{\theta_{\text{GP}}^{\text{attend}}})\right],$ $\quad H_{t}=-\sum_{j=1}^{V}p_{t,j}\log p_{t,j},$ (3.1)
$\text{where}\quad\mathbf{p}_{t}=\pi_{\theta}(\cdot\mid\mathcal{R}_{<t},x;T)$ $=\text{Softmax}\left(\frac{\mathbf{z}_{t}}{\tau}\right).$
Here, $\mathcal{J}_{\text{GRPO}}$ is the PPO-clip loss term from GRPO in Eq. 2.1, $V$ is the vocabulary size, $\mathbf{z}_{t}\in\mathbb{R}^{V}$ are the pre-softmax logits, and $\tau$ is the decoding temperature. Critically, the entropy bonus coefficient $\gamma_{s}$ is not a fixed hyperparameter but is dynamically set based on the curriculum level $s$ of the current sample. For $s=1$, we set $\gamma_{\text{easy}}\approx 0$, as we do not need to explicitly reward exploration when the agent should be learning to exploit reliable specialist knowledge. For $0<s<1$, we use a moderate positive bonus $\gamma_{\text{medium}}>0$ to encourage policy diversity and prevent the agent from becoming overconfident in the face of conflicting information. For $s=0$, we apply a strong positive bonus $\gamma_{\text{hard}}\gg\gamma_{\text{medium}}$ to aggressively incentivize exploration, compelling the model to break from the misleading specialist consensus.
@@
Input: Dataset $\mathcal{D}=\{x_{v}^{(i)},x_{t}^{(i)},y^{*(i)}\}_{i=1}^{N}$, policy model $\pi_{\theta}$, old policy $\pi_{\text{old}}$, group size $G$, specialist responses $y_{d}^{(i)}$.